OpenStudy (naga447):

Can someone please show me how to find the answer? What is the exact distance from (-5, 1) to (3, 0)

OpenStudy (naga447):

square root of sixty-nine units square root of sixty-five units square root of sixty-three units square root of fifty units

OpenStudy (anonymous):

\[\LARGE \bf d=\sqrt{(y_1-y_2)^{2} +(x_1-x_2)^{2}}\]\[\LARGE \bf d=\sqrt{(1-0)^{2} +(-5-3)^{2}}\]\[\LARGE \bf d=\sqrt{1^{2} +(-8)^{2}}\]\[\LARGE \bf d=\sqrt{1+64}\]\[\LARGE \bf d=\sqrt{65}\]

OpenStudy (anonymous):

(you can't simplify √65 )

OpenStudy (naga447):

Okay thanks

OpenStudy (naga447):

If only I knew what that means....

OpenStudy (anonymous):

You welcome... but you don't know what, what means ?

OpenStudy (naga447):

I don't know what your doing to get the answer...

OpenStudy (anonymous):

I am using the distance formula.